# Energy extraction through Comisso-Asenjo process from a Kerr-Newman black hole in perfect fluid dark matter

### v1

@@ Abstract
In this work, we provide a thorough analysis of energy extraction via magnetic reconnection, a novel mechanism recently proposed by Comisso and Asenjo, for a Kerr-Newman black hole immersed in a perfect fluid dark matter (PFDM) background. Our studies focus on the impact of black hole spin $a$, electric charge $Q$ and PFDM parameter $\lambda$ on the horizons, ecoregions and circular geodesics at the equatorial plane of this black hole, and how they further influence the reconnection efficiency and energy extraction rate. Our results show that the size of ergoregion does not vary monotonically with increasing dark matter parameters $\lambda$, but it can significantly increase at faster spins ($a>0.8$) as the dark matter parameter $\lambda$ decreases, given the electric charge stays within $Q\in[0.2,0.5]$. We identify the optimal conditions for the combination of $a$, $Q$ and $\lambda$ that enable efficient energy extraction even when the black hole is not rapidly spinning. The Kerr-Newman black hole in PFDM allows for achieving high energy extraction rates comparable to those of most previously studied rotating black holes, which typically require near-extremal spin to reach similar efficiency levels.
pacs:
11.25.Hf, 04.60.-m, 04.70.-s
I INTRODUCTION
@@
Astrophysical black holes, among the most enigmatic and dense objects predicted by Einstein’s general theory of relativity, have captivated scientists for their profound influence on high-energy astrophysical phenomena. In recent years, the development of advanced instruments such as gravitational wave interferometers Abbott et al. (2016, 2019) and the Event Horizon Telescope (EHT) Akiyama et al. (2019a, b, 2022a, 2022b) have significantly enhanced the study of black hole physics, opening new avenues for phenomenological research. One of the primary focuses in high-energy astrophysics is to identify the sources of relativistic jets and beams such as gamma-ray bursts (GRBs). Black holes, as natural candidates capable of driving energy production in the universe, are suggested responsible for producing relativistic jets originating from two main sources: the gravitational potential energy of matter falling toward a black hole during accretion, and the electromagnetic field energy generated in the vicinity of a black hole Misner (1972); Penrose (1969); Penrose and Floyd (1971). This area of research continues to be highly valued and actively pursued in the quest to understand how these energetic astrophysical phenomena are connected to the mechanics of black hole energy extraction.
@@
The Penrose process is one of the first well-known frameworks to address energy extraction from a rotating black hole Penrose (1969). It involves a particle entering the ergosphere of a rotating black hole, and splitting into two parts. A space-like Killing vector inside the ergosphere allows one part to fall into the black hole with negative energy relative to an observer at infinity, while the other escapes with more energy than the original particle, resulting in a decrease in the black hole’s total energy and mass. For a rotating black hole with mass $M$ and dimensionless spin $a$, the maximum amount of energy that can be extracted is its rotational energy $E_{tot}=(M-M_{irr})c^{2}$, where $M_{irr}=M\sqrt{(1+\sqrt{1-a^{2}})/2}$ is the irreducible mass Christodoulou (1970) and $c$ is the speed of light. For a maximally rotating black hole ($a=1$), it could potentially lose up to 29% of its mass-energy equivalence (0.29 $Mc^{2}$). Although theoretically appealing, the Penrose process is less practical in most astrophysical contexts due to the extreme velocities ($>0.5c$) required for the particles involved and its insufficiency in extracting significant rotational energy Bardeen et al. (1972); Wald (1974). Building on Penrose’s foundational work, researchers have proposed several alternative energy extraction mechanisms, including the collisional Penrose process Piran et al. , superradiant scattering Teukolsky and Press (1974), the Blandford-Znajek (BZ) process Blandford and Znajek (1977), and the magnetohydrodynamic Penrose process Takahashi et al. (1990). Among these, the BZ process has gained wide acceptance for explaining the energetic jets observed in active galactic nuclei (AGNs) McKinney and Gammie (2004); Hawley and Krolik (2006); Komissarov and McKinney (2007); Tchekhovskoy et al. (2011) and GRBs Lee et al. (2000); Tchekhovskoy et al. (2008); Komissarov and Barkov (2009). According to the BZ process, a spinning zero-charge black hole in a uniform magnetic field generated by the accretion disk induces an electric field that directs charged particles. The charged plasma medium allows a charged current to form and twist magnetic field lines into a helix, draining energy and powering jets.
Recently, Comisso and Asenjo extended a novel mechanism for energy extraction via magnetic reconnection within the ergosphere of a rotating black hole Comisso and Asenjo (2021), denoted as the Comisso-Asenjo (CA) process, to distinguish from an earlier exploration in a similar avenue by Kiode and Arai Koide and Arai (2008). Magnetic reconnection is commonly observed in solar flares and the earth’s magnetosphere, where antiparallel magnetic field lines break and reconnect. This mechanism produces thin current sheets and plasmoids that facilitate the rapid conversion of magnetic field energy into kinetic energy carried by released plasma (e.g. Comisso et al., 2016, 2017; Daughton et al., 2009; Uzdensky and Loureiro, 2016; Bhattacharjee et al., 2009). Similar to the Penrose process, magnetic reconnection generates accelerated plasma outflows with equal velocities in opposite directions; one outflow is driven against the black hole’s spin and the other escapes, enabling the energy extraction from the black hole.
General-relativistic kinetic simulations (e.g. Ripperda et al., 2020; Parfrey et al., 2019; Komissarov, 2005; East and Yang, 2018) suggest reconnection process commonly and repeatedly occurs due to the frame-dragging effect in the magnetosphere of rapidly rotating black holes. Multi-frequency radio telescope observations have verified the presence of strong magnetic fields around supermassive black holes, which can influence the dynamics of accretion Eatough et al. (2013). The EHT collaboration also observed polarized emission in the magnetosphere of the supermassive black hole in M87* Akiyama et al. (2019b). In Comisso and Asenjo’s seminal work, they demonstrated that the fast plasma motion facilitated by the reconnection process can extract energy much more efficiently, especially for black holes with high spins and strong magnetic fields, and can potentially surpass the energy extraction efficiency during the BZ process. Inspired by the original study of the CA process in Kerr black holes, recent research on various parametrized rotating black holes Zhang (2024a); Chen et al. (2024); Zhang (2024b); Shaymatov et al. (2023); Khodadi et al. (2023); Ye et al. (2023); Li and Yuan (2023); Li et al. (2023); Wang et al. (2022); Carleo et al. (2022); Liu (2022); Khodadi (2022) has revealed differing impacts on energy extraction and plasma energization efficiencies.
Another critical aspect that has gained attention but has not been thoroughly studied yet is the role of dark matter in the vicinity of a rotating charged black hole during energy extraction via reconnection. Indirect observations suggest that dark matter is abundant and everywhere in the observable universe, constituting at least approximately 27% of the total matter Rubin et al. (1980); Persic et al. (1996); Bertone and Hooper (2018). Dark matter is assumed to form halos around black holes in most theories and contributes to the effective black hole mass Jusufi et al. (2019); Kiselev (2005); Navarro et al. (1996); Kiselev (2003). Studying black hole models in such realistic astrophysical backgrounds is crucial as dark matter can affect geodesics, leading to significant consequences for the dynamics and energetics of these systems. Recently, black hole solutions in the presence of dark matter have been studied using a perfect fluid dark matter (PFDM) model Rahaman et al. (2011), setting the stage for calculating black hole shadow Hou et al. (2018), geodesics Das et al. (2021); Xu et al. (2018), thermodynamics Liang et al. (2023); Rakhimova et al. (2023) and superradiance Liu et al. (2024).
In this paper, we explore energy extraction via magnetic reconnection for a Kerr-Newman black hole in PFDM using the framework pioneered by Comisso and Ansenjo. Specifically, we investigate how the presence of dark matter affects the efficiency and dynamics of the energy extraction process. By incorporating PFDM in the black hole model, we can gain more accurate insights into the interplay between black holes and their realistic surroundings and a deeper understanding of high-energy astrophysics and cosmology.
@@
The paper is organized as follows. In section II, we briefly introduce the Kerr-Newman solution in the presence of PFDM. In section III, we show the impact of black hole spin $a$, electric charge $Q$, and the PFDM parameter $\lambda$ on the event horizons, ergoregion and circular geodesics of the black hole concerned. In section IV, we introduce and analyze the energy extraction via the CA mechanism in detail by studying the parameter and phase space and computing the extraction rate and reconnection efficiency. We summarize our results in section V. We have assumed geometrized units $c=G=1$ throughout this work.
@@
II KERR-NEWMAN BLACK HOLE IN PERFECT FLUID DARK MATTER
@@
We start with the action for Einstein’s gravity theory minimally coupled to a U(1) gauge field immersed in perfect fluid dark matter (PFDM), which takes the following form Das et al. (2021)
@@
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=8\pi G\left\{T^{M}_{\mu\nu}+T^{DM}_{\mu\nu}\right\}=8\pi GT_{\mu\nu},$ (2)
where $T^{M}_{\mu\nu}$ is the energy-momentum tensor due to the electromagnetic fields generated by the electric charge $Q$
@@
where the lapse function $f(r)$ is solved by requiring $\epsilon=\frac{3}{2}$
@@
To complete the Kerr-Newman solution in PFDM, the black hole spin $a$ is incorporated using the Newman-Janis algorithm (readers can refer to Das et al. (2021) for more details). In Boyer-Lindquist coordinates $(t,r,\theta,\phi)$, where $r$, $\theta$ and $\phi$ represent the radial distance, the polar angle and the azimuthal angle, respectively, the final form of the metric can be described as
@@
$g_{t\phi}=-\frac{2a\sin^{2}\theta}{\rho^{2}}\left[2Mr-Q^{2}-\lambda r\ln\left(\frac{r}{|\lambda|}\right)\right],g_{\theta\theta}=\rho^{2}$ (10)
@@
III GEODESICS OF KERR-NEWMAN BLACK HOLE IN PFDM
Since energy extraction occurs in the ergosphere region of a rotating black hole, it is crucial to first investigate the horizons and stationary limits of a Kerr-Newman black hole in PFDM. Unlike a regular Kerr black hole, where the inner boundary of the ergosphere coincides with the outer event horizon, the structures of horizons and stationary limit surfaces in this context are more complex and interesting. They depend sensitively on the three parameters in the metric: black hole spin $a$, electric charge $Q$, and the PFDM parameter $\lambda$. In the CA mechanism, it is assumed that magnetic reconnection that occurs in the bulk plasma rotating circularly around the black hole is observed at infinity. Thus we solve for the inner ($r_{h-})$ and outer ($r_{h+}$) boundaries of horizons specifically at the equatorial plane $(\theta=\frac{\theta}{2})$ by requiring
@@ III GEODESICS OF KERR-NEWMAN BLACK HOLE IN PFDM
It should be noted here that the maximum allowed value for $\lambda$ is $2M$ for $\lambda>0$ and the minimum allowed value for $\lambda$ is $-7.18M$ when $\lambda<0$ Xu et al. (2018). Without losing generality in methodology, we exclusively work with the positive values of $\lambda$ throughout the paper. Differing from the restriction of $M^{2}\geq a^{2}+Q^{2}$ for the regular Kerr-Newman black hole, the presence of a dark matter background counteracts the negative gravitational effect caused by the electric charge. Consequently, larger values of $Q$ are permissible in this context, provided that $\lambda$ is sufficiently large to allow the logarithmic term in $\Delta$ to turn negative. While it is impossible to obtain analytical expressions for $r_{h\pm}$ and $r_{L\pm}$ due to the complexity introduced by the logarithmic term in $\Delta$, we perform a series of numerical studies on Eqn 14 and Eqn 15 with various combinations of $(a,Q,\lambda)$.
As illustrated in Fig 1, we observe several trends of how $(a,Q,\lambda)$ could affect the positions of the two horizons of this black hole, leading to significant changes in the size of the ergoregion. Panels (a) and (b) in Fig. 1 show an example of a standard peanut-shaped ergoregion, bounded by the orange (outer horizon) and red (outer boundary of ergosphere) contours for two $\lambda$ values with $a$ and $Q$ fixed. A greater dark matter component could significantly increase the outer horizon while simultaneously decreasing the inner horizon, resulting in a shrinking ergoregion. In fact, we believe this variation is caused by the combination of both electric charge and dark matter background, compared to the minimal impact of $\lambda$ on the inner horizon of the rotating non-charged black hole in the PFDM case discussed in Liu et al. (2024).
Panel 1(c) shows the outer horizon $r_{h+}$ as a function of the dark matter parameter $\lambda$ for three values of electric charge when the spin is fixed at $a=0.5$. Interestingly, the outer horizon initially decreases as $\lambda$ increases, then begins to rapidly increase with $\lambda$ once $\lambda$ surpasses a critical value $\lambda_{c}$ located at the minimum of $r_{h+}$. This critical value $\lambda_{c}$, interpreted as the point of reflection, has also been noted in previous studies Das et al. (2021) and has been observed for the shadow of the rotating black hole in PFDM Hou et al. (2018).
@@
Since non-spinning test particles can only stably orbit the black hole beyond or at the innermost stable circular orbit (ISCO), one favored condition for energy extraction via reconnection is that ISCO orbit must reside at or within the outer boundary of the ergosphere. Therefore, we are also interested in examining how the electric charge and dark matter background impact the circular geodesics of the Kerr-Newman black hole in PFDM. We start by considering a test particle with Lagrangian
@@
The radial equation of motion in the case of null geodesics can then be obtained from the conservation of the Hamiltonian $\mathcal{H}=p_{\mu}\dot{x}^{\mu}-\mathcal{L}$, yielding:
$\dot{r}^{2}=\frac{1}{4}\left[\left(E(r^{2}+a^{2})-al_{z}\right)^{2}-\Delta(l_{z}-aE)^{2}\right].$ (18)
And the effective potential $V_{eff}$ for the Kerr-Newman black hole in PFDM is given by Das et al. (2021)
$\begin{split}V_{eff}=-\left(1+g_{tt}\dot{t}^{2}+2g_{t\phi}\dot{t}\dot{\phi}+g_{\phi\phi}\dot{\phi}^{2}\right)\\ ~{}~{}=-\frac{2M}{r^{3}}(aE-l_{z})^{2}+\frac{Q^{2}}{r^{4}}(aE-l_{z})^{2}\\ +\frac{\lambda}{r^{3}}\ln\left(\frac{r}{|\lambda|}\right)-\frac{1}{r^{2}}(a^{2}E^{2}-l_{z}^{2})+\frac{\Delta}{r^{2}}.\end{split}$ (19)
Figure 2: Innermost Stable Circular Orbits ($r_{isco}$) as a function of the black hole spin $a$ for the black hole charge $Q\in\{0.3,0.5,0.7\}$ with fixed $\lambda=1.2$.
To determine circular rotating photon orbits ($r_{p\pm}$), we impose the requirements $\dot{r}=0$ and $\partial_{r}V_{eff}=0$, where the $\pm$ signs represent the co-rotating and the counter-rotating photon orbits, respectively. To find the innermost stable circular orbits ($r_{isco}$), we further require the condition $\partial^{2}_{r}V_{eff}=0$. As shown in Fig 2, $r_{isco}$ decreases slowly with black hole spin $a$ in lower values of $a$ and steeply as $a$ approaches higher values. Also, the range of $r_{isco}$ increases as the electric charge increases, but it always resides within the outer boundary of the ergosphere.
IV ENERGY EXTRACTION BY CAMISSO-ASENJO PROCESS
To conveniently analyze the plasma properties during the energy extraction process by magnetic reconnection, we consider a locally nonrotating frame known as the Zero Angular Momentum Observer (ZAMO) frame Bardeen et al. (1972), which we will denote by hats for the rest of the paper. The spacetime in the ZAMO frame is locally flat (Minkowski) and its squared line element is given by

### v2

# Energy extraction through magnetic reconnection from a Kerr-Newman black hole in perfect fluid dark matter

@@ Abstract
In this work, we provide a thorough analysis of energy extraction via magnetic reconnection, a novel mechanism recently proposed by Comisso and Asenjo, for a Kerr-Newman black hole immersed in a perfect fluid dark matter (PFDM) background. Our studies focus on the impact of black hole spin $a$, electric charge $Q$ and PFDM parameter $\lambda$ on the horizons, ergoregion and circular geodesics at the equatorial plane of this black hole, and how they further influence the reconnection efficiency and energy extraction rate. Our results show that the outer horizon and the size of ergoregion do not vary monotonically with increasing dark matter parameters $\lambda$ until reaching its critical value $\lambda_{c}$ due to the combined counteracting effect between the black hole’s charge and dark matter parameter. We identify the optimal combinations of $a$, $Q$ and $\lambda$ that allow for efficient energy extraction and high extracted power, even when the black hole is not spinning near its extremal limit. Our results ease the stringent conditions observed in other rotating black holes, where achieving comparable levels of extracted power and reconnection efficiency typically requires a near-extremal spin.
keywords:
Dark matter , Black holes , Rotating charged , Geodesics , Energy extraction , Magnetic reconnection
[…]
[CfA]organization=Harvard-Smithsonian Center for Astrophysics, city=Cambridge MA, postcode=02138, country=USA
1 Introduction
@@
Astrophysical black holes, among the most enigmatic and dense objects predicted by Einstein’s general theory of relativity, have captivated scientists for their profound influence on high-energy astrophysical phenomena. In recent years, the development of advanced instruments such as gravitational wave interferometers [1, 2] and the Event Horizon Telescope (EHT) [3, 4, 5, 6] have significantly enhanced the study of black hole physics, opening new avenues for phenomenological research. One of the primary focuses in high-energy astrophysics is to identify the sources of relativistic jets and beams such as gamma-ray bursts (GRBs). Black holes, as natural candidates capable of driving energy production in the universe, are suggested to be responsible for producing relativistic jets originating from two main sources: the gravitational potential energy of matter falling toward a black hole during accretion, and the electromagnetic field energy generated in the vicinity of a black hole [7, 8, 9]. This area of research continues to be highly valued and actively pursued in the quest to understand how these energetic astrophysical phenomena are connected to the mechanics of black hole energy extraction.
@@
The Penrose process is one of the first well-known frameworks to address energy extraction from a rotating black hole [8]. It involves a particle entering the ergosphere of a rotating black hole, and splitting into two parts. A space-like Killing vector inside the ergosphere allows one part to fall into the black hole with negative energy relative to an observer at infinity, while the other escapes with more energy than the original particle, resulting in a decrease in the black hole’s total energy and mass. For a rotating black hole with mass $M$ and dimensionless spin $a$, the maximum amount of energy that can be extracted is its rotational energy $E_{tot}=(M-M_{irr})c^{2}$, where $M_{irr}=M\sqrt{(1+\sqrt{1-a^{2}})/2}$ is the irreducible mass [10] and $c$ is the speed of light. For a maximally rotating black hole ($a=1$), it could potentially lose up to 29% of its mass-energy equivalence (0.29 $Mc^{2}$). Although theoretically appealing, the Penrose process is less practical in most astrophysical contexts due to the extreme velocities ($>0.5c$) required for the particles involved and its insufficiency in extracting significant rotational energy [11, 12]. Building on Penrose’s foundational work, researchers have proposed several alternative energy extraction mechanisms, including the collisional Penrose process [13], superradiant scattering [14], the Blandford-Znajek (BZ) process [15], and the magnetohydrodynamic Penrose process [16]. Among these, the BZ process has gained wide acceptance for explaining the energetic jets observed in active galactic nuclei (AGNs) [17, 18, 19, 20] and GRBs [21, 22, 23]. According to the BZ process, a rotating, electrically neutral black hole in a surrounding magnetic field generated by the accretion disk, induces an electric field that accelerates charged particles in the surrounding plasma. This interaction generates a current, twisting the magnetic field lines into a helical structure, resulting in rotational energy extraction from the black hole that is channeled into powering relativistic jets.
In addition to large-scale powerful jets, frequent bursts of flares originating near the event horizon of accreting black holes have been observed across multiple wavelengths. These include gamma-ray flares in the GeV range from active galactic nuclei (AGN) [24, 25, 26, 27], energetic TeV flares from M87* [28, 29], and powerful X-ray and infrared flares from Sgr A* [30, 31, 32, 33]. However, the exact mechanism responsible for triggering these bright flares in the vicinity of black holes remains debated. Koide and Arai [34] were the first to explore magnetic reconnection as a possible energy extraction mechanism from a rotating black hole. Magnetic reconnection is commonly observed in solar flares and the earth’s magnetosphere, where antiparallel magnetic field lines break and reconnect. This mechanism produces thin current sheets and plasmoids that facilitate the rapid conversion of magnetic field energy into kinetic energy carried by released plasma [e.g. 35, 36, 37, 38, 39]. Similar to the Penrose process, magnetic reconnection generates accelerated plasma outflows with equal velocities in opposite directions; one outflow is driven against the black hole’s spin and the other escapes, enabling the energy extraction from the black hole. Multi-frequency radio telescope observations have verified the presence of strong magnetic fields around supermassive black holes, which can influence the dynamics of accretion [40]. The EHT collaboration also observed polarized emission in the magnetosphere of the supermassive black hole in M87* [4], confirming twisting magnetic field lines near the horizon that possibly initiate reconnection. General-relativistic MHD simulations [e.g. 41, 42, 43, 44, 45] suggest that the reconnection process commonly and repeatedly occurs due to the frame-dragging effect in the magnetosphere of rapidly rotating black holes.
Recently, Comisso and Asenjo extended this mechanism by requiring fast magnetic reconnection within the ergosphere of a rotating black hole and computed the energy extraction rate and reconnection efficiency [46]. They demonstrated that the fast plasma motion facilitated by the reconnection process can extract energy much more efficiently, especially for black holes with high spins and strong magnetic fields, and can potentially surpass the energy extraction efficiency during the BZ process. To distinguish their approach from the earlier work by Koide and Arai, we refer to this as the Comisso-Asenjo (CA) process. Numerous investigations have been conducted for energy extraction via the CA process in a variety of scenarios, such as for Kerr-de Sitter black holes [47], rotating non-Kerr, rotating hairy or regular black holes [48, 49, 50], rotating magnetized black holes [51], rotating black holes with broken Lorentz symmetry [52], Kerr black holes in modified gravity [53], Konoplya-Rezzolla-Zhidenko parametrized black holes [54], rotating black holes with a U(1) charge [55] or a string charge [56], spinning wormholes [57] and Kerr plunging region [58]. These studies have revealed differing impacts on energy extraction and plasma energizationf efficiencies by the relevant parameters in their black hole models. However, another critical parameter that has not yet been thoroughly studied is the role of dark matter in the vicinity of a rotating charged black hole during energy extraction via reconnection, as dark matter is also known to have significant impacts on the black hole’s spacetime [59, 60]. Indirect observations and the Standard Model of cosmology both suggest that dark matter is abundant and everywhere in the observable universe, constituting at least approximately 27% of the total matter [61, 62, 63]. Dark matter is assumed to form halos around black holes in most theories and contributes to the effective black hole mass [64, 65, 66, 67]. Studying black hole models in such realistic astrophysical backgrounds is crucial as dark matter can affect geodesics, leading to significant consequences for the dynamics and energetics of these systems. A series of models to study the interaction between dark matter and black holes have been proposed. One dark matter model type is based on an astrophysical density profile of the dark matter halo that depends on the density parameter and characteristic radius, as dark matter appears to form a density structure near the black hole but disappears asymptotically. This approach is commonly used in galaxy formation and structure simulations, such as the popular NFW, Burkert and Einasto profiles for studying the dark matter halo structure in the cold dark matter (CDM) model [66, 68, 69, 70, 71, 72]. Another alternative dark matter model is to describe dark matter as a perfect fluid [73, 74, 75], which has the advantage of providing a well-defined stress-energy tensor leading to an analytical solution, making it easier to study how dark matter interacts gravitationally with black holes, including how it accretes onto the black hole or affects spacetime curvature. Many black hole phenomenological studies have been conducted using the perfect fluid dark matter (PFDM) approach, such as calculating black hole shadow [59, 60], thin accretion disks [76], geodesics [77, 78, 79], thermodynamics [80, 81] and superradiance [82].
@@
In this paper, we explore energy extraction via magnetic reconnection for a Kerr-Newman black hole in a PFDM background following the framework pioneered by Comisso and Ansenjo. Specifically, we investigate how the presence of dark matter affects the efficiency and dynamics of the energy extraction process. The paper is organized as follows. In section 2, we briefly introduce the Kerr-Newman solution in the presence of PFDM. In section 3, we show the impact of black hole spin $a$, electric charge $Q$, and the PFDM parameter $\lambda$ on the event horizons, ergoregion and circular geodesics of the black hole concerned. In section 4, we introduce and analyze the energy extraction via the CA mechanism in detail by studying the parameter and phase space and computing the extraction rate and reconnection efficiency. We summarize our results in section 5. We have assumed geometrized units $c=G=1$ throughout this work.
@@
2 KERR-NEWMAN BLACK HOLE IN PERFECT FLUID DARK MATTER
@@
Since Kiselev first formulated spherically symmetric black hole metrics in the presence of PFDM [73], subsequent developments have extended these solutions to include rotation, leading to the Kerr-(A)dS solution [78] and the Kerr-Newman solution [77]. To review the construction of a rotating charged black hole solution in PFDM, we start with the action for Einstein’s gravity theory minimally coupled to a U(1) gauge field immersed in perfect fluid dark matter (PFDM), which takes the following form
@@
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=8\pi G\left\{T^{M}_{\mu\nu}+T^{DM}_{\mu\nu}\right\}=8\pi GT^{s}_{\mu\nu},$ (2)
where $T^{s}_{\mu\nu}$ represents the total energy-momentum tensor sourcing the black hole in consideration. $T^{M}_{\mu\nu}$ is the energy-momentum tensor due to the electromagnetic fields generated by the electric charge $Q$
@@
where the lapse function $f(r)$ is solved for by requiring $\epsilon=\frac{3}{2}$
@@
To complete the Kerr-Newman solution in PFDM, the black hole spin $a$ is incorporated using the Newman-Janis algorithm. In Boyer-Lindquist coordinates $(t,r,\theta,\phi)$, where $r$, $\theta$ and $\phi$ represent the radial distance, the polar angle and the azimuthal angle, respectively, the final form of the metric can be described as [77]
@@
$g_{t\phi}=-\frac{a\sin^{2}\theta}{\rho^{2}}\left[2Mr-Q^{2}-\lambda r\ln\left(\frac{r}{|\lambda|}\right)\right],g_{\theta\theta}=\rho^{2}$ (10)
@@
This solution reduces to a Kerr-Newman black hole without the presence of PFDM background when the parameter $\lambda=0$.
3 GEODESICS OF KERR-NEWMAN BLACK HOLE IN PFDM
Since energy extraction occurs in the ergosphere region of a rotating black hole, it is crucial to first investigate the horizons and stationary limits of a Kerr-Newman black hole in PFDM. Unlike a regular Kerr black hole, where the inner boundary of the ergosphere coincides with the outer event horizon, the structures of horizons and stationary limit surfaces in this context are more complex and interesting. They depend sensitively on the three parameters in the metric: black hole spin $a$, electric charge $Q$, and the PFDM parameter $\lambda$. In the CA mechanism, it is assumed that magnetic reconnection that occurs in the bulk plasma rotating circularly around the black hole is observed at infinity. Thus we solve for the inner ($r_{h-})$ and outer ($r_{h+}$) boundaries of horizons specifically at the equatorial plane $(\theta=\frac{\pi}{2})$ by requiring
@@ 3 GEODESICS OF KERR-NEWMAN BLACK HOLE IN PFDM
It should be noted here that both positive and negative values of $\lambda$ are permitted and the horizon structure dependence on $\lambda$ with its limits has been thoroughly investigated in [79] for the case of rotating non-charged black hole in PFDM. Here, we provide an extensive study on the horizon when the black hole’s charge $Q$ is also present. Without losing generality in methodology, we exclusively work with the positive values of $\lambda$ throughout the paper. Differing from the restriction of $M^{2}\geq a^{2}+Q^{2}$ for the regular Kerr-Newman black hole, the presence of a dark matter background counteracts the negative gravitational effect caused by the electric charge. Consequently, larger values of $Q$ are permissible in this context, provided that $\lambda$ is sufficiently large to allow the logarithmic term in $\Delta$ to turn negative. While it is impossible to obtain analytical expressions for $r_{h\pm}$ and $r_{L\pm}$ due to the complexity introduced by the logarithmic term in $\Delta$, we perform a series of numerical studies on Eqn 14 and Eqn 15 with various combinations of $(a,Q,\lambda)$.
As illustrated in Fig 1, we observe several trends of how $(a,Q,\lambda)$ could affect the positions of the two horizons of this black hole, leading to significant changes in the size of the ergoregion. Panels (a) and (b) in Fig. 1 show an example of a standard peanut-shaped ergoregion, bounded by the orange (outer horizon) and red (outer boundary of ergosphere) contours for two $\lambda$ values with $a$ and $Q$ fixed. A greater dark matter component could significantly increase the outer horizon, resulting in a shrinking ergoregion.
Panel 1(c) shows the outer horizon $r_{h+}$ as a function of the dark matter parameter $\lambda$ for three values of electric charge when the spin is fixed at $a=0.5$. Interestingly, the outer horizon initially decreases as $\lambda$ increases, then begins to rapidly increase with $\lambda$ once $\lambda$ surpasses a critical value $\lambda_{c}$ located at the minimum of $r_{h+}$. This critical value $\lambda_{c}$, interpreted as the point of reflection, has also been noted in previous studies [77] and has been observed for the shadow of the rotating black hole in PFDM [59].
@@
Magnetic reconnection occurs within the bulk plasma orbiting circularly around the black hole, with circular orbits extending from infinity down to the limiting circular photon orbit. We now turn our attention to examine how the electric charge and dark matter background impact the circular null geodesics of the Kerr-Newman black hole in PFDM. We start by considering a test particle with Lagrangian
@@
The radial equation of motion can then be obtained from the conservation of the Hamiltonian $\mathcal{H}=p_{\mu}\dot{x}^{\mu}-\mathcal{L}$, yielding
$\dot{r}^{2}=\frac{1}{r^{4}}\left[\left(E(r^{2}+a^{2})-al_{z}\right)^{2}-\Delta(l_{z}-aE)^{2}\right]+\frac{\Delta}{r^{2}}\epsilon,$ (18)
where $\epsilon=-1,0,1$ for timelike, null and spacelike particles, respectively. In the case of null geodesics, the radial equation becomes
$\begin{split}\dot{r}^{2}=\frac{1}{r^{4}}\left[\left(E(r^{2}+a^{2})-al_{z}\right)^{2}-\Delta(l_{z}-aE)^{2}\right]=f(r)\end{split}$ (19)
Figure 2: Panel (a) illustrates the co-rotating photon orbit as a function of dark matter parameter $\lambda$ with fixed black hole charge $Q=0.3$ and three spin values $a=0.2,0.8,1.0$. Panel (b) illustrates the co-rotating photon orbit as a function of black hole charge $Q$ with fixed dark matter parameter $\lambda=1.2$ and three spin values $a=0.2,0.8,1.0$.
[…]
$\begin{split}r_{p}^{2}+\frac{2M}{r_{p}}(a-D)^{2}-\frac{Q^{2}}{r_{p}^{2}}(a-D)^{2}-\frac{\lambda}{r_{p}}\ln\left(\frac{r_{p}}{|\lambda|}\right)(a-D)^{2}+(a^{2}-D^{2})&=0,\\ 2r_{p}-\frac{2M}{r_{p}^{2}}(a-D)^{2}+\frac{2Q^{2}}{r_{p}^{3}}(a-D)^{2}+\frac{\lambda}{r_{p}^{2}}\ln\left(\frac{r_{p}}{|\lambda|}\right)(a-D)^{2}-\frac{\lambda}{r_{p}^{2}}(a-D)^{2}&=0.\end{split}$ (20)
Solving for $D$ and subtituting it into the first equation above yields the following equation for the co-rotating photon orbit $r_{p}$ (here we focus on co-rotating orbit for magnetic reconnection occurring inside the ergosphere)
$\begin{split}6Mr_{p}^{3}-4Q^{2}r_{p}^{3}-3\lambda r_{p}^{3}\ln\left(\frac{r_{p}}{|\lambda|}\right)+\lambda r_{p}^{3}-2r_{p}^{4}+2a\sqrt{2r_{p}^{4}\left(2Mr_{p}-2Q^{2}-\lambda r_{p}\ln\left(\frac{r_{p}}{|\lambda|}\right)+\lambda r_{p}\right)}=0\end{split}$ (21)
Since obtaining an analytical expression for the $r_{p}$ is not feasible, we report numerical results for the co-rotating photon orbit. As shown in panel (a) of Fig 2, the photon orbit $r_{p}$ exhibits a similar trend to the outer horizon in its dependence on the dark matter parameter $\lambda$ while keeping charge $Q$ fixed- it initially decreases with increasing $\lambda$ until reaching a critical value $\lambda_{c}$. Meanwhile, panel (b) of Fig 2 shows that $r_{p}$ decreases monotonically with increasing value of charge $Q$ while keeping dark matter parameter $\lambda$ fixed. These trends also align with the fact that when dark matter interacts with black holes, it effectively increases the mass, acting as an attractive gravitational charge (when $\lambda>\lambda_{c}$) that mitigates the repulsive effect of the electric charge. The results from this section clearly demonstrate the strong dependence of the Kerr-Newman black hole in PFDM spacetime on the black hole charge $Q$ and the dark matter parameter $\lambda$. Therefore, it is compelling to investigate their influences on the rate and efficiency of energy extraction via magnetic reconnection, which we explore in the next section.
4 ENERGY EXTRACTION BY CAMISSO-ASENJO PROCESS
To investigate the mechanism behind the energetic bursts of flares caused by charged particles near a black hole, we examine the energy extraction process via magnetic reconnection, as recently proposed by Comisso and Asenjo [46]. Specifically, we focus on the strong frame-dragging effect in a rotating charged black hole immersed in PFDM and analyze the efficiency of its rotational energy being extracted by the surrounding plasma in the accretion disk. First, we consider a locally nonrotating frame known as the Zero Angular Momentum Observer (ZAMO) frame [11], which we will denote by hats for the rest of the paper. The spacetime in the ZAMO frame is locally flat (Minkowski) and its squared line element is given by